# Can a Quantum Computer Simulate Nuclear Magnetic Resonance Spectra Better than a Classical One?

### v1

@@ A.2 Computing the Spectral Function from the Correlation Function
We can interpret these formulas (26) and (27) for the magnetic moments as correlation functions, respectively denoted $C_{YY}$ and $C_{XY}$, making use of the notation
$C_{\alpha\beta}\left(t\right)=\text{Tr}\left[\left(M^{\alpha}\right)^{\dagger}e^{-iHt}M^{\beta}e^{+iHt}\right]$ (29)
@@
$\begin{split}C_{++}\left(t\right)&=\text{Tr}\left[\left(M^{+}\right)^{\dagger}e^{-iHt}M^{+}e^{+iHt}\right]\\ &=\text{Tr}\left[M^{-}e^{-iHt}M^{+}e^{+iHt}\right],\end{split}$ (30)
@@
If we further make use of the rotational symmetry of the Hamiltonian under the transformation
$X\to Y~;~Y\to-X,$ (32)
@@
$\begin{split}C_{++}\left(t\right)&=\sum_{n,m}\langle E_{n}|M^{-}e^{-iHt}|E_{m}\rangle\langle E_{m}|M^{+}e^{+iHt}|E_{n}\rangle\\ &=\sum_{n,m}e^{-i\left(E_{m}-E_{n}\right)t}\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle\end{split}$ (34)
@@
$C_{++}\left(w\right)=\int_{0}^{+\infty}C_{++}\left(t\right)e^{-i\omega t}e^{-\eta t}d\omega$ (35)
@@
$\int_{0}^{+\infty}e^{-i\left(\omega+E_{m}-E_{n}\right)t}e^{-\eta t}d\omega=\frac{-i}{\omega+E_{m}-E_{n}-i\eta}$ (36)
@@
$C_{++}\left(w\right)=-i\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\omega+E_{m}-E_{n}-i\eta}.$ (37)
Taking the real part of this expression, we obtain
$\text{Re}\left[C_{++}\left(w\right)\right]=\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega+E_{m}-E_{n}]^{2}},$ (38)
@@ Appendix C Parameter Source
In Table 1, we list the molecules which we study in our analysis, and indicate the source of the chemical shifts and couplings which appear in the NMR Hamiltonian. All of the active nuclei for the molecules mentioned below are 1H nuclei, except for the case of Diphosphane, which in addition to containing 20 1H nuclei, also possesses two 31P nuclei.
@@ Appendix D Additional Convergence Data
In this Appendix we present some additional convergence data for exactly solvable systems. Figures 13 through 18 display the results for selected molecules of size 8, 10, and 12, for the usual parameter regimes. The convergence data is again generally what we would expect, although there are perhaps some stronger finite-size effects. We also have data for two 9-site molecules, t-Butyl chloride and 2-Methyl-2-cyanopropane, but both of these immediately converge to within machine precision for even the smallest cluster sizes, for every possible parameter regime.

### v2

@@ A.2 Computing the Spectral Function from the Correlation Function
We can interpret these formulas (30) and (31) for the magnetic moments as correlation functions, respectively denoted $C_{XY}$ and $C_{YY}$, making use of the notation
$C_{\alpha\beta}\left(t\right)=\text{Tr}\left[\left(\hat{M}^{\alpha}\right)^{\dagger}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{\beta}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\right]$ (33)
@@
$\begin{split}C_{++}\left(t\right)&=\text{Tr}\left[\left(\hat{M}^{+}\right)^{\dagger}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{+}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\right]\\ &=\text{Tr}\left[\hat{M}^{-}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{+}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\right],\end{split}$ (34)
@@
If we further make use of the cyclic nature of the trace, together with the rotational symmetry of the Hamiltonian under rotations around the Z axis,
$R\hat{M}^{x}R^{\dagger}=\hat{M}^{y};~R\hat{M}^{y}R^{\dagger}=-\hat{M}^{x};~R\mathrm{e}^{\pm i\hat{\cal{H}}_{R}t}R^{\dagger}=\mathrm{e}^{\pm i\hat{\cal{H}}_{R}t}$ (36)
@@
$\begin{split}C_{++}\left(t\right)&=\sum_{n,m}\langle E_{n}|\hat{M}^{-}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}|E_{n}\rangle\\ &=\sum_{n,m}\mathrm{e}^{-i\left(E_{m}-E_{n}\right)t}\langle E_{n}|\hat{M}^{-}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}|E_{n}\rangle\end{split}$ (38)
@@
$C_{++}\left(w\right)=\int_{0}^{+\infty}C_{++}\left(t\right)\mathrm{e}^{-i\omega t}\mathrm{e}^{-\eta t}dt$ (39)
@@
$\int_{0}^{+\infty}\mathrm{e}^{-i\left(\omega+E_{m}-E_{n}\right)t}\mathrm{e}^{-\eta t}dt=\frac{-i}{\omega+E_{m}-E_{n}-i\eta}$ (40)
@@
$C_{++}\left(w\right)=-i\sum_{n,m}\frac{\langle E_{n}|\hat{M}^{-}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}|E_{n}\rangle}{\omega+E_{m}-E_{n}-i\eta}.$ (41)
Since $\hat{M}^{-}$ and $\hat{M}^{+}$ are conjugates of each other, the numerator of this expression is real, and so taking the real part of the full expression, we obtain
$\text{Re}\left[C_{++}\left(w\right)\right]=\eta\sum_{n,m}\frac{\langle E_{n}|\hat{M}^{-}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}|E_{n}\rangle}{\eta^{2}+[\omega+E_{m}-E_{n}]^{2}},$ (42)
@@
We note that the eigenstates above are those of $\hat{\cal{H}}_{R}$, the Hamiltonian in the rotating frame (indeed, our solver performs the diagonalization on the rotating frame Hamiltonian). However, the only effect this has on the spectral function is to shift all of the frequencies by the Larmor frequency (as one would intuitively expect), and since we ultimately plot everything in terms of the chemical shift anyways, this is ultimately a relatively unimportant distinction. We also emphasize that while it is not necessary to include nuclei other than hydrogen in the $\hat{M}^{\pm}$ operators, it is still crucial to include them in the Hamiltonian, since this will generate important splittings in the hydrogen spectrum.
@@ Appendix C Parameter Source
In Table 1, we list the molecules which we study in our analysis, and indicate the source of the chemical shifts and couplings which appear in the NMR Hamiltonian. All of the active nuclei for the molecules mentioned below are 1H nuclei, except for the case of a.) Diphosphane, which in addition to containing 20 1H nuclei, also possesses two 31P nuclei, as well as the case of b.) Triphenylphosphine oxide, which in addition to containing 15 1H nuclei, also possesses one 31P nucleus.
@@ Appendix D Additional Convergence Data
In this Appendix we present some additional convergence data for exactly solvable systems. Figures 17 through 22 display the results for selected molecules of size 8, 10, and 12, for the usual parameter regimes. The convergence data is again generally what we would expect, although there are perhaps some stronger finite-size effects. We also have data for two 9-site molecules, t-Butyl chloride and 2-Methyl-2-cyanopropane, but both of these immediately converge to within machine precision for even the smallest cluster sizes, for every possible parameter regime.